# references.bib

### v2

@@ Acknowledgement
Figure 7(a) shows how the computational efficiency varies with the switching thresholds, $\Omega$, for the SIRS model with demography presented in the previous section, and compares it to the Doob-Gillespie method. We see that as we increase the switching threshold, the Jump-Switch-Flow method requires more computational time to simulate the scenario for four years, but is always more efficient than the Doob-Gillespie method.
@@
Figure S7: 7(a) The running time to simulate the scenario for four years for an initial population size of $N(0)=10^{5}$, with varying switching threshold, $\Omega$ (grey to purple), compared with the Doob-Gillespie (‘G’) method (yellow). The endemic steady state number of infectious individuals is also presented (vertical dashed line). 7(b) The running time for the Jump-Switch-Flow method for an appropriate threshold choice is approximately constant as the population size varies (for example $\Omega=10^{3}$). If a poor switching threshold is chosen, the running time begins to scale exponentially (for example $\Omega=10^{5}$). For the Tau-Leaping method, the running time is constant, and then begins to grow as the rate parameter increases with population size. For the Doob-Gillespie method, the running time scales exponentially with the population size.
To observe how the computational efficiency of our Jump-Switch-Flow method varies with the total (population) size of the system, we fix the switching threshold at $\Omega=10^{3}$, as this value is sufficient to capture the majority of the key dynamics we are interested in. Figure 7(b) compares the Jump-Switch-Flow method (purple) to the Doob-Gillespie method (yellow) and to the computationally more efficient approximation Tau-Leaping method (red), noting the log-log scale. We can see that the computational efficiency of the Doob-Gillespie method scales exponentially with system size, as expected. The Tau-Leaping method initially has constant computational time scaling, up until $N(0)=10^{4}$, where it starts to grow exponentially with the system size. This change in time scaling is caused by the Tau-Leaping method computing Poisson random numbers differently, depending on the rate parameter. For sufficiently large rate parameters, this random number generation increases in computational complexity, leading to the exponential computational time scaling with system size. In contrast, our Jump-Switch-Flow method is constant in time up to an initial system size of $N(0)=10^{5}$, after which it begins to slowly decrease (for $\Omega=10^{1}$ and $\Omega=10^{3}$). This decrease is caused by more of the scenario being simulated in the flowing state, as the system becomes highly deterministic with a larger population.
@@
S3.5 Simulation experiments: Inference with simulated data, parameter estimates
@@
To test how the Jump-Switch-Flow method behaves when used to perform parameter estimation, we first generated synthetic data of a known SIRS mode with demography. We simulate data for 400 days using the parameter values given in Figure 2C, with the Doob-Gillespie algorithm, and select a trajectory which experiences epidemic fade-out (at day 250), and an initial condition of two infected individuals ($I(0)=2$), no recovered individuals ($R(0)=0$), and a total population of $N(0)=10^{5}$ individuals. We track the number of infectious individuals within the population each day, and use a truncated binomial distribution to add noise into the data set, reflecting a realistic measurement process. We then couple our Jump-Switch-Flow method, using a switching threshold of $\Omega=10^{3}$, with the open-source particle filter package pypfilt (Moss2024), using 2,000 particles. We first fix the demographic parameters to $\kappa=\mu=1/(85\times 365)$ (since these are assumed to be standard within a given population). We perform the inference on the data set for the first 100 days of the epidemic to estimate $\beta$, $\gamma$ and $\omega$, and then use our parameter estimates to predict forward in time the possible trajectories based on the obtained predictions. Using these predictions, we also estimate the probability of epidemic fade-out occurring at times $t=100$ to $t=400$ days.
@@
Figure S8 shows the posterior distributions of parameters. The dotted black lines show the true value used to produce the simulated data, and the dashed red lines indicate the initial bounds on the prior estimates. We see that both $\beta$ and $\gamma$ are well estimated, with the posteriors centered around the true values. However $\omega$ not well estimated. This is because $\omega$ acts on the order of $1$ year, while the window of data used to fit the parameters to the data is 100 days ($\approx 0.3$ years).
@@
Figure S8: Posterior distributions of parameters for the SIRS model with demography, obtained via a particle filter with 2,000 particles, and the Jump-Switch-Flow sampler. The red dashed lines indicate the initial bounds on prior samples. The vertical dotted black lines indicate the true value used to produce the simulated data.
@@
S4 Inference Study: Estimating TEIRV model parameters using a particle filter
@@
We implemented the Refractory Cell Model described by the authors for the viral load data obtained from nasal swabs, using our Jump-Switch-Flow method with a switching threshold on all compartments $\Omega=10^{2}$. We implemented this in particle filter pypfilt Moss2024 to estimate the model parameters, using 6,000 particles per simulation. In previous analysis ke2022daily, the decay of virion was fixed at $c=10$ virions/day, and the eclipse of cells to become infectious as $k=4$ cells/day, leaving the remaining parameters to be estimated. Therefore, we similarly fix $c=10$ virions/day and $k=4$ cells/day.
S4.1 Parameter Estimation and Viral Clearance Prediction
@@
Figure 9 of the main text highlights how the estimated viral loads closely resemble the data collected from the 6 patients. We also show how the estimated $R_{0}$ values for the patients vary. These $R_{0}$ estimates are consistent with the original study by ke2022daily. However, understanding how the parameter estimates vary between patients assists in understanding how the fitting process behaves.
@@
The posterior distributions for each of the 6 patient’s parameter estimates are provided in Figure S9. These posterior distributions are found by fitting a violin envelope from the 6,000 estimates obtained from the particle filtering process. We also show the prior distributions, as red dashed lines, which show the lower and upper bounds for the uniform initial guesses.
@@
Figure S9: Posterior distributions of the model parameters and initial viral load for each of the selected SARS-CoV-2 infections. The red dashed lines indicate the initial bounds on prior samples.

### v3

@@ Acknowledgement
Figure 10(a) shows how the computational efficiency varies with the switching thresholds, $\Omega$, for the SIRS model with demography presented in the previous section, and compares it to the Doob-Gillespie method. We see that as we increase the switching threshold, the Jump-Switch-Flow method requires more computational time to simulate the scenario for four years, but is always more efficient than the Doob-Gillespie method. This increased efficiency is because the Jump-Switch-Flow method utilises the Next Reaction Method to compute the stochastic events, a method known to be more computationally more efficient than the direct, first reaction, Doob-Gillespie method.
@@
Figure SI10: 10(a) The running time to simulate the scenario for four years for an initial population size of $N(0)=10^{5}$, with varying switching threshold, $\Omega$ (grey to purple), compared with the Doob-Gillespie (‘G’) method (yellow). The endemic steady state number of infectious individuals (vertical dashed line), susceptible individuals (vertical solid line), and recovered individuals (vertical dotted line) are presented. 10(b) The running time for the Jump-Switch-Flow method for an appropriate threshold choice is approximately constant as the population size varies (for example $\Omega=10^{3}$). If a poor switching threshold is chosen, the running time begins to scale exponentially (for example $\Omega=10^{5}$). For the Tau-Leaping method, the running time is constant, as we use fixed step size of $\Delta\tau=0.01$ (solid line) and $\Delta\tau=0.1$ (dashed line). We also present the Tau-hybrid method, in both Python and C++, with various different time meshes, C++ with a fine mesh ($\delta t=0.01$, black solid line) and a corse mesh ($\delta t=0.1$, black dashed line), and Python with a corse mesh ($\delta t=0.1$, grey dotted line).
To observe how the computational efficiency of our Jump-Switch-Flow method varies with the total (population) size of the system, we consider three switching thresholds: $\Omega=10^{1}$ (dotted purple), $\Omega=10^{3}$ (dashed purple), and $\Omega=10^{5}$ (solid purple). Figure 10(b) compares the Jump-Switch-Flow method (purple) to the Doob-Gillespie method (yellow), to the computationally more efficient approximation Tau-Leaping method ($\Delta\tau=0.01$, dashed red, and $\Delta\tau=0.1$, solid red), and to the Tau-hybrid method ($\Delta t=0.01$ solid black, and $\Delta t=0.1$ dashed black) provided by GillesPy2, noting the log-log scale. We simulate the SIRS system described above, for a total time of 4×3654$\times$ 3654 × 365 days. We can see that the computational efficiency of the Doob-Gillespie method scales exponentially with system size, as expected. The Tau-Leaping method initially has constant computational time scaling. In contrast, our Jump-Switch-Flow method is constant in time up to an initial system size of $N(0)=10^{5}$, after which it begins to slowly decrease (for $\Omega=10^{1}$ and $\Omega=10^{3}$). This decrease is caused by more of the scenario being simulated in the flowing state, as the system becomes highly deterministic with a larger population. We also observe our Jump-Switch-Flow method increasing exponentially between $N(0)=10^{4}$, for the largest switching threshold of $\Omega=10^{5}$ (solid purple). We observe our Jump-Switch-Flow method performing better, in comparison to Doob-Gillespie exact method, as we use the Next Reaction Method to sample the Jump times. This is despite our implementation being required to perform extra computation to check the regime of the Jump-Switch-Flow process, and implement any relevant regime changes (even though they are not required in this particular instance). We also observe that for large, realistic population sizes (i.e. $N(0)\geq 10^{5}$), the Tau-hybrid method with the corse time mesh (black dashed line) performs at best comparable to, or worse than the Jump-Switch-Flow method, if a suitable $\Omega$ is chosen (such as $\Omega=10^{3}$. Moreover, as we specify a finer time mesh (solid black line) we observe that the Tau-hybrid method performs worse across all population sizes. We anticipate this is due to the Tau-hybrid method utilising an arbitrary ODE solver, an adaptive $\Delta\tau$ that is calculated at each step, and also due to the dynamic switching behaviour between stochastic and deterministic regimes that requires extra computation of the propensities to determine if a switching event should occur.
@@
SI4.5 Simulation experiments: Inference with simulated data, parameter estimates
@@
To test how the Jump-Switch-Flow method behaves when used to perform parameter estimation, we first generated synthetic data of a known SIRS mode with demography. We simulate data for 400 days using the parameter values given in Figure 2C, with the Doob-Gillespie algorithm, and select a trajectory which experiences epidemic fade-out (at day 250), and an initial condition of two infected individuals ($I(0)=2$), no recovered individuals ($R(0)=0$), and a total population of $N(0)=10^{5}$ individuals. We track the number of infectious individuals within the population each day, and use a truncated binomial distribution to add noise into the data set, reflecting a realistic measurement process. We then couple our Jump-Switch-Flow method, using a switching threshold of $\Omega=10^{3}$, with the open-source particle filter package pypfilt [35], using 2,000 particles. We first fix the demographic parameters to $\kappa=\mu=1/(85\times 365)$ (since these are assumed to be standard within a given population). We perform the inference on the data set for the first 100 days of the epidemic to estimate $\beta$, $\gamma$ and $\omega$, and then use our parameter estimates to predict forward in time the possible trajectories based on the obtained predictions. Using these predictions, we also estimate the probability of epidemic fade-out occurring at times $t=100$ to $t=400$ days.
@@
Figure SI11 shows the posterior distributions of parameters. The dotted black lines show the true value used to produce the simulated data, and the dashed red lines indicate the initial bounds on the prior estimates. We see that both $\beta$ and $\gamma$ are well estimated, with the posteriors centered around the true values. However $\omega$ not well estimated. This is because $\omega$ acts on the order of 111 year, while the window of data used to fit the parameters to the data is 100 days (0.3$\approx$ 0.3≈ 0.3 years).
@@
Figure SI11: Posterior distributions of parameters for the SIRS model with demography, obtained via a particle filter with 2,000 particles, and the Jump-Switch-Flow sampler. The red dashed lines indicate the initial bounds on prior samples. The vertical dotted black lines indicate the true value used to produce the simulated data.
@@
SI5 Inference Study: Estimating TEIRV model parameters using a particle filter
@@
We implemented the Refractory Cell Model described by the authors for the viral load data obtained from nasal swabs, using our Jump-Switch-Flow method with a switching threshold on all compartments $\Omega=10^{2}$. We implemented this in particle filter pypfilt [35] to estimate the model parameters, using 6,000 particles per simulation. In previous analysis [23], the decay of virion was fixed at $c=10$ virions/day, and the eclipse of cells to become infectious as $k=4$ cells/day, leaving the remaining parameters to be estimated. Therefore, we similarly fix $c=10$ virions/day and $k=4$ cells/day. The prior distributions used for the analysis with the TEIRV model are given in Table SI5.
Parameter Prior
[…]
$c$ Fixed at 10.010.010.0
$k$ Fixed at 4.04.04.0
[…]
Table SI5: The prior distributions used for the analysis with the TEIRV model.
SI5.1 Parameter Estimation and Viral Clearance Prediction
@@
Figure 9 of the main text highlights how the estimated viral loads closely resemble the data collected from the 6 patients. We also show how the estimated $R_{0}$ values for the patients vary. These $R_{0}$ estimates are consistent with the original study by [23]. However, understanding how the parameter estimates vary between patients assists in understanding how the fitting process behaves.
@@
The posterior distributions for each of the 6 patient’s parameter estimates are provided in Figure SI12. These posterior distributions are found by fitting a violin envelope from the 6,000 estimates obtained from the particle filtering process. We also show the prior distributions, as red dashed lines, which show the lower and upper bounds for the uniform initial guesses.
@@
Figure SI12: Posterior distributions of the model parameters and initial viral load for each of the selected SARS-CoV-2 infections. The red dashed lines indicate the initial (uniform) bounds on prior samples.